A 528 Hz sound wave travels through a substance. The wavelength of the sound wave is measured to be 1.9 meters. What is the speed of sound in the substance?
Answer #1

speed=frequency * wavelength

speed = 528 * 1.9 = 1003.2 m/s.
Answer #2

time period=i/528

speed=1.9*528=1,003.2 m/s
Answer #3

v = f*\lambda

v = 528*1.9

v = 1003.2 m/s